# Black Hole Orbit Classification: The Synergistic Effects of Cloud Strings, Gauss-Bonnet Terms, and Non-Commutative Parameters in Identifying WGC candidate Models: WGC as WCCC protector

### v1

@@ IV.1 $CaseI:\alpha<0$
$\begin{split}H=\frac{\sqrt{4+\frac{\left(1-\sqrt{1+4\left(\frac{4\left(\arctan\left(\frac{r}{\sqrt{\Xi\pi}}\right)-\frac{\sqrt{\Xi\pi}\,r}{\Xi\pi+r^{2}}\right)m}{\pi r^{3}}-\frac{q}{r^{4}}+\frac{a}{r^{2}}\right)\alpha}\right)r^{2}}{\alpha}}}{2\sin\!\left(\theta\right)r}.\end{split}$ (23)
@@
$\begin{split}\phi^{r}=\frac{\csc\!\left(\theta\right)\left(\varphi_{1}+\varphi_{2}+\left(\sqrt{\Xi}\,r^{2}\left(a\,r^{2}-2q\right)\pi^{\frac{3}{2}}-10m\,r^{2}\pi\Xi-6m\pi^{2}\Xi^{2}\right)r^{2}\right)}{2\sqrt{\Xi}\,\varphi_{3}\pi\left(\Xi\pi+r^{2}\right)^{2}r^{4}}.\end{split}$ (24)
@@
Figure 7: Fig (7a): Metric function with $a=0.6,\alpha=-0.1$ and different $\Xi$ for Black Hole with Cloud of strings, Fig (7b): The normal vector field $n$ in the $(r-\theta)$ plane. The photon sphere is located at $(r,\theta)=(1.800667715288,1.57)$ with respect to $(a=0.6,q=0.6,\Xi=0.001,\alpha=-0.1,m=1)$, (7c): The photon spheres are located at $(r,\theta)=(0.78056312953,1.57)$ and $(r,\theta)=(1.125202804843,1.57)$ with respect to $(a=1.4,q=0.6,\Xi=0.0155,\alpha=-0.1,m=1)$ for naked singularity with CLOUD OF STRINGS
@@
Table 3: *Unauthorized region: is the region where the roots of $\phi$ equations become negative or imaginary. $R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@
$A=E^{2}r^{2}-L^{2}\left(1+\frac{\left(1-\sqrt{1+4\left(\frac{4\left(\arctan\left(\frac{r}{\sqrt{\Xi\pi}}\right)-\frac{\sqrt{\Xi\pi}\,r}{\Xi\pi+r^{2}}\right)m}{\pi r^{3}}-\frac{q}{r^{4}}+\frac{a}{r^{2}}\right)\alpha}\right)r^{2}}{4\alpha}\right).$ (26)
$B=r^{2}\left(1+\frac{\left(1-\sqrt{1+4\left(\frac{4\left(\arctan\left(\frac{r}{\sqrt{\Xi\pi}}\right)-\frac{\sqrt{\Xi\pi}\,r}{\Xi\pi+r^{2}}\right)m}{\pi r^{3}}-\frac{q}{r^{4}}+\frac{a}{r^{2}}\right)\alpha}\right)r^{2}}{4\alpha}\right).$ (27)
@@
Figure 8: Fig (8a): $\beta$ diagram for structure in black hole form, (8b): $\beta$ diagram for structure in naked singularity form
@@
Table 4: *Unauthorized region: is the region where the roots of $\phi$ equations become negative or imaginary. $R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@
Figure 9: Fig (9a): $\beta$ diagram for structure in the black hole form, (9b): $\beta$ diagram for structure in the naked singularity form
@@
Table 5: *Unauthorized region: is the region where the roots of $\phi$ equations become negative or imaginary. $R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@
Figure 10: Fig (10a): $\beta$ diagram for structure in the black hole form, (10b): $\beta$ diagram for structure in the naked singularity form
@@ VII Appendix A: NON-COMMUTATIVE SCHWARZSCHILD BLACK HOLE
Table 7: *Unauthorized region: is the region where the roots of $\phi$ equations become negative or imaginary. $R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@ VIII Appendix B: NON-COMMUTATIVE 4D Einstein-Gauss-Bonnet
Table 8: *Unauthorized region: is the region where the roots of $\phi$ equations become negative or imaginary. $R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.

### v2

@@ IV.1 $CaseI:\alpha<0$
$\begin{split}\mathbb{H}=\frac{\sqrt{4+\frac{\left(1-\sqrt{1+4\left(\frac{4\left(\arctan\left(\frac{r}{\sqrt{\Xi\pi}}\right)-\frac{\sqrt{\Xi\pi}\,r}{\Xi\pi+r^{2}}\right)m}{\pi r^{3}}-\frac{q}{r^{4}}+\frac{a}{r^{2}}\right)\alpha}\right)r^{2}}{\alpha}}}{2\sin\!\left(\theta\right)r}.\end{split}$ (23)
@@
$\begin{split}\varphi_{r}=\frac{\csc\!\left(\theta\right)\left(\varphi_{1}+\varphi_{2}+\left(\sqrt{\Xi}\,r^{2}\left(a\,r^{2}-2q\right)\pi^{\frac{3}{2}}-10m\,r^{2}\pi\Xi-6m\pi^{2}\Xi^{2}\right)r^{2}\right)}{2\sqrt{\Xi}\,\varphi_{3}\pi\left(\Xi\pi+r^{2}\right)^{2}r^{4}}.\end{split}$ (24)
@@
Figure 7: Fig (7a): Metric function with $a=0.6,\alpha=-0.1$ and different $\Xi$ for Black Hole with Cloud of strings, Fig (7b): The photon sphere is location at $(r,\theta)=(1.800667715288,1.57)$ with respect to $(a=0.6,q=0.6,\Xi=0.001,\alpha=-0.1,m=1)$, (7c): The photon spheres location at $(r,\theta)=(0.78056312953,1.57)$ and $(r,\theta)=(1.125202804843,1.57)$ with respect to $(a=1.4,q=0.6,\Xi=0.0155,\alpha=-0.1,m=1)$ for naked singularity with CLOUD OF STRINGS
@@
Table 3: *Unauthorized region: The region with negative or imaginary roots of $\varphi$.
$R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@
$\mathbb{A}=E^{2}r^{2}-L^{2}\left(1+\frac{\left(1-\sqrt{1+4\left(\frac{4\left(\arctan\left(\frac{r}{\sqrt{\Xi\pi}}\right)-\frac{\sqrt{\Xi\pi}\,r}{\Xi\pi+r^{2}}\right)m}{\pi r^{3}}-\frac{q}{r^{4}}+\frac{a}{r^{2}}\right)\alpha}\right)r^{2}}{4\alpha}\right).$ (26)
$\mathbb{B}=r^{2}\left(1+\frac{\left(1-\sqrt{1+4\left(\frac{4\left(\arctan\left(\frac{r}{\sqrt{\Xi\pi}}\right)-\frac{\sqrt{\Xi\pi}\,r}{\Xi\pi+r^{2}}\right)m}{\pi r^{3}}-\frac{q}{r^{4}}+\frac{a}{r^{2}}\right)\alpha}\right)r^{2}}{4\alpha}\right).$ (27)
@@
Figure 8: Fig (8a): With respect to $a=0.6,q=0.6,\Xi=0.001,\alpha=-0.1,m=1$, the $\beta$ diagram in the black hole form and (8b): With respect to $a=0.6,q=0.6,\Xi=0.0155,\alpha=-0.1,m=1$, the $\beta$ diagram in the naked singularity mode
@@
Table 4: *Unauthorized region: The region with negative or imaginary roots of $\varphi$
$R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@
Figure 9: Fig (9a): With respect to $a=0.8,q=0.6,\Xi=0.001,\alpha=0.01,m=1$, the $\beta$ diagram in the black hole form and (9b): With respect to $a=0.8,q=0.6,\Xi=0.004,\alpha=0.01,m=1$, the $\beta$ diagram in the naked singularity mode
@@
Table 5: *Unauthorized region: The region with negative or imaginary roots of $\varphi$.
$R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@
Figure 10: Fig (10a): With respect to $a=0.6,q=1,\Xi=0.001,\alpha=0.3,m=3$, the $\beta$ diagram in the black hole form and (10b): With respect to $a=0.6,q=1,\Xi=0.15,\alpha=0.3,m=3$, the $\beta$ diagram in the naked singularity mode
@@ VII Appendix A: NON-COMMUTATIVE SCHWARZSCHILD BLACK HOLE
Table 7: *Unauthorized region: The region with negative or imaginary roots of $\varphi$.
$R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@ VIII Appendix B: NON-COMMUTATIVE 4D Einstein-Gauss-Bonnet
Table 8: *Unauthorized region: The region with negative or imaginary roots of $\varphi$.
$R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.